OpenStudy (anonymous):

Derek kicks a soccer ball off the ground and in the air with an initial velocity of 31 feet per second. Using the formula H(t) = −16t2 + vt + s, what is the maximum height the soccer ball reaches? 14.2 feet 14.6 feet 15.0 feet 15.3 feet

OpenStudy (anonymous):

i think it's 15.0 feet

hero (hero):

How do you know this? Did you do any calculations to figure that out?

OpenStudy (anonymous):

yes

hero (hero):

Show me what you've done.

OpenStudy (anonymous):

H(t) = -16t^2 + 31t = -t(16t - 31) At the maximum height, t = (1/2)(0+31/16) = .96875 H( .96875) = 15.0 ft

hero (hero):

What happened to the + s part of the formula?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

H(t) = −16t^2 + vt + s dH(t)/dt = −32t + v 0 = −32t + 31 t = 31/32

OpenStudy (anonymous):

substitute into H(t) = −16t^2 + vt + s

OpenStudy (anonymous):

s is unknown

hero (hero):

If s is not known, you need to figure out what its value is. You can't just ignore it.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

well thats where i got stuck on